# DKDS: A Benchmark Dataset of Degraded Kuzushiji Documents with Seals for Detection and Binarization

### v1

@@ Abstract
Kuzushiji, a pre-modern Japanese cursive script, can currently be read and understood by only a few thousand trained experts in Japan. With the rapid development of deep learning, researchers have begun applying Optical Character Recognition (OCR) techniques to transcribe Kuzushiji into modern Japanese. Although existing OCR methods perform well on clean pre-modern Japanese documents written in Kuzushiji, they often fail to consider various types of noise, such as document degradation and seals, which significantly affect recognition accuracy. To the best of our knowledge, no existing dataset specifically addresses these challenges. To address this gap, we introduce the Degraded Kuzushiji Documents with Seals (DKDS) dataset as a new benchmark for related tasks. We describe the dataset construction process, which required the assistance of a trained Kuzushiji expert, and define two benchmark tracks: (1) text and seal detection and (2) document binarization. For the text and seal detection track, we provide baseline results using multiple versions of the You Only Look Once (YOLO) models for detecting Kuzushiji characters and seals. For the document binarization track, we present baseline results from traditional binarization algorithms, traditional algorithms combined with K-means clustering, and Generative Adversarial Network (GAN)–based methods. The DKDS dataset and the implementation code for baseline methods are available at https://ruiyangju.github.io/DKDS.
@@
Kuzushiji, Seals, Pre-modern Japanese Documents, Benchmark Dataset, Object Detection, Document Binarization
@@ 1 Introduction
Pre-modern Japanese documents were mainly written in Kuzushiji111くずし字, a form of Japanese cursive script (草書体)., a pre-modern Japanese cursive script widely used in official records, personal letters, and literary works. Kuzushiji is characterized by complex, flowing strokes and highly variable forms, which differ markedly from modern Japanese characters [ueki2021survey]. Educational reforms in the early 20th century introduced standardized textbooks and led to reduced instruction in Kuzushiji. Following the adoption of modern kana orthography in 1946, most contemporary native Japanese speakers are unable to read Kuzushiji texts without specialized training [hashimoto2017Kuzushiji].
With the remarkable advances in deep learning for computer vision and natural language processing, researchers have increasingly applied neural network models to Kuzushiji character recognition [le2019human, le2020automated]. While trained models typically achieve satisfactory performance in recognizing Kuzushiji characters seen during training, their performance often drops significantly when tested on unseen (zero-shot) Kuzushiji characters. For instance, the framework proposed by Ishikawa et al. [ishikawa2024japanese] achieved an accuracy of only approximately 48% in such zero-shot scenarios. This result presents the critical need to construct high-quality Kuzushiji datasets, including special cases, to improve model generalization and capabilities in complex recognition tasks.
The primary Kuzushiji datasets include Kuzushiji-MNIST, Kuzushiji-49, and Kuzushiji-Kanji [^2] Kanji refers to Chinese-origin characters used in Japanese. [clanuwat2018deep]. The Kaggle competitions organized using these datasets has further advanced both research and practical applications [kitamoto2019progress, kitamoto2020kaggle]. Together, these datasets and competitions cover a wide range of Kuzushiji characters and provide a standardized evaluation platform for researchers. With these resources, researchers can perform not only Kuzushiji character classification but also related tasks such as Kuzushiji detection and generation [soan2023], thereby advancing the development of Kuzushiji character recognition techniques and enhancing model performance.
Several methods have been proposed for the Kuzushiji character recognition task. Fujita et al. [fujita2024learning] employed Transformer [vaswani2017attention] models to analyze and recognize Kuzushiji characters, leveraging attention mechanisms to capture relationships between adjacent characters and accurately separate character regions and boundaries. Ngo et al. [ngo2021recurrent] introduced RNN-Transducer [graves2012sequence] models for handwritten Kuzushiji character recognition, effectively exploiting both visual and text information from the input images and achieving satisfactory performance.
@@
In addition, KuroNet [clanuwat2019kuronet, lamb2020kuronet] is an end-to-end model based on a residual U-Net [ronneberger2015u] architecture, designed to capture long-range context, process extensive vocabularies, and accommodate non-standard character layouts, thereby enabling the recognition of pre-modern Japanese documents. On the Kuzushiji dataset provided by the Kaggle competitions [kitamoto2019progress, kitamoto2020kaggle], KuroNet achieved an accuracy exceeding 90%. To facilitate the deployment of OCR models on mobile devices, Clanuwat et al. [clanuwat2021miwo] developed the “miwo” application, which allows users to upload or photograph Kuzushiji documents and automatically transcribe the characters into modern Japanese.
@@
Among the various types of noise, the overlap between Kuzushiji characters and seals poses a major challenge. Serving as symbols of ownership and collector identity, seals were widely used in ancient and pre-modern Asian document collections [li2021character]. They are typically inscribed with stylized ancient scripts and may contain not only the collector’s name but also phrases expressing personal wishes, interests, or social status. In ancient and pre-modern documents, seals usually appear as red marks that frequently overlap with characters, which can blur the characters and lead to OCR errors.
To address this challenge, we introduce the Degraded Kuzushiji Documents with Seals (DKDS) dataset, developed in collaboration with a trained Kuzushiji expert. This novel benchmark dataset comprises degraded pre-modern Japanese documents written in Kuzushiji, with various types of seals randomly added to simulate common types of interference. As shown in Fig. 2, we define two task tracks for the DKDS dataset: (1) text and seal detection, and (2) document binarization. The former track aims to accurately detect text (i.e., Kuzushiji characters) and seal regions within the documents, while the latter focuses on preserving the primary textual content while removing noise, including stains and seals, thereby providing clean input for subsequent OCR tasks.
@@
    We introduce a novel benchmark dataset, named Degraded Kuzushiji Documents with Seals (DKDS), developed in collaboration with a trained Kuzushiji expert. To the best of our knowledge, this is the first publicly available dataset specifically addressing the challenge of Kuzushiji characters overlapped with seals.
@@
    We define the text and seal detection task, which serves as a preliminary step for recognizing Kuzushiji characters and transcribing them into modern Japanese. Baseline results for this task are also presented using multiple versions of the YOLO models.
@@
    We define the document binarization task specifically for pre-modern Japanese document images. Compared to existing binarization datasets, the DKDS dataset presents additional challenges due to the overlap of Kuzushiji characters with seals. Baseline results for this task are provided, including traditional binarization algorithms, traditional algorithms combined with K-means clustering, and Generative Adversarial Network (GAN)–based methods.
@@ 2.1 Related Dataset
The datasets relevant to this work can be categorized into three domains: (1) Kuzushiji document datasets, (2) degraded pre-modern document binarization datasets, and (3) seal datasets.
Existing research on pre-modern Japanese documents written in Kuzushiji [clanuwat2018deep, le2019human] lacks a dataset that specifically addresses the challenge of overlaps between characters and seals, which can significantly reduce recognition accuracy.
Most publicly available datasets for degraded ancient document binarization focus on documents written in Latin-based alphabetic systems. For instance, the documents in the Document Image Binarization Contest (DIBCO) series datasets [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2012icfhr, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019] are primarily written in English, while the Bickley Diary (BD) dataset [deng2010binarizationshop], Synchromedia Multispectral Ancient Document Images (SMADI) [hedjam2013historical], and Persian Heritage Image Binarization Dataset (PHIBD) [ayatollahi2013persian] all contain documents written in European languages. In contrast, no publicly available document binarization datasets exist for pre-modern Japanese documents, particularly those written in Kuzushiji.
Furthermore, no existing degraded document datasets include seals as the primary source of interference. Therefore, we aim to randomly add seals to the collected pre-modern Japanese documents written in Kuzushiji. However, existing publicly available seal datasets are all unsuitable, as the seals they contain are either not ancient Asian seals, suffer from poor image quality, or have complex backgrounds. Specifically, the Chinese-Seal Dataset (CSD) [xu2025chinese] focuses on modern Chinese seals; DDI-100 [zharikov2020ddi] contains Western (English-language) seal samples; and SPODS [nandedkar2016spods] comprises seal images from contemporary office settings. Although MiikeMineStamps [buitrago2021miikeminestamps] includes some Japanese seal images, they are generally low-resolution and feature non-uniform backgrounds, making background removal difficult.
@@ 2.2 Text and Seal Detection
The diverse shapes of ancient seals, their curved inscriptions, and the frequent occurrence of text overlap pose significant challenges for seal processing and text analysis and recognition. As a critical preliminary step for Kuzushiji character recognition, text detection is particularly important [li2023yolo]. Similarly, seal detection is important for subsequent seal processing.
@@
However, the scarcity of ancient seal samples and the prevalence of low-quality or degraded images severely limit available training data and annotations for related tasks [li2021character]. For instance, Micenkov et al. [micenkov2011stamp] proposed an automatic seal segmentation system for insurance documents that achieved 84% accuracy. Yu et al. [yu2023icdar] developed a modern seal dataset for official and financial contexts, introducing two tasks: seal title text detection and end-to-end seal title recognition. To the best of our knowledge, no dataset currently exists for text and seal detection in degraded pre-modern Japanese Kuzushiji documents.
@@
The YOLO series of models [jocher2023yolov8, wang2024yolov9, wang2024yolov10, jocher2024yolo11] achieve an excellent balance between detection accuracy and model size, demonstrating outstanding performance across a wide range of object detection tasks [ju2023fracture, chien2025yolov8, chien2024yolov9, ju2024yolov8, ju2024global]. Researchers have explored the application of the YOLO series of models in Kuzushiji character detection [li2023yolo] and seal detection [bento2025performance]. For instance, Bento et al. [bento2025performance] merged and adapted the StaVer [micenkov2011stamp] and DDI-100 [zharikov2020ddi] datasets to evaluate the performance of YOLOv8, YOLOv9, YOLOv10, and YOLO11 on their newly constructed dataset for seal detection. Similarly, Sun et al. [sun2024ra] enhanced the YOLOv8 architecture and proposed RA-YOLOv8, achieving higher detection accuracy. Nevertheless, due to the lack of seal data based on Kuzushiji documents, these improved YOLO models cannot be directly applied to this specific scenario.
@@
2.3 Document Binarization
Document binarization is an important preprocessing step for OCR and layout analysis, effectively removing various types of noise from document images, such as paper yellowing, text fading, ink bleeding, stains, and seals. The Document Image Binarization Contest (DIBCO), one of the most popular competitions in this field, has provided benchmark datasets from 2009 to 2019 [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2012icfhr, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019]. Researchers have proposed a range of neural network models evaluated on these datasets, including those based on GANs [suh2022two, ju2023ccdwt, ju2024mfe], Transformers [souibgui2022docentr], and diffusion models [saifullah2023coldbin, yang2023docdiff], which have significantly advanced the field. However, most existing datasets contain ancient and pre-modern document images based on Latin alphabets, while few include Kuzushiji documents, particularly those containing overlapping seals and text.
Figure 4: Examples of our collected imperial seals from the Qing dynasty, with backgrounds removed.
Figure 5: The overall workflow of the proposed DKDS dataset construction includes raw data collection, detection annotations, initial binarization ground-truth generation, verification, and manual correction. The initial binarization ground-truth was generated using the model proposed by Ju et al. [ju2024three], while the verification was conducted by a trained Kuzushiji expert.
@@ 3.1 Raw Data Collection
The raw Kuzushiji document data were obtained from the Center for Open Data in the Humanities (CODH) [genjimonogatari]. Specifically, we selected the dataset corresponding to book ID “200003803”, titled Genji Monogatari (The Tale of Genji) [^3] https://codh.rois.ac.jp/char-shape/book/200003803/. Examples from this book are shown in Fig 3. This dataset was chosen for two main reasons: (1) its historical significance and widespread popularity; and (2) the availability of detailed OCR annotations provided by CODH, including 237 unique character types and a total of 11,132 characters. These data can be used for training and evaluating OCR models in future research.
Furthermore, to date, there is no publicly available high-resolution dataset of ancient Japanese seals. To address this gap, we collected 219 seals from three emperors of the Qing dynasty (i.e., Kangxi (1661–1722), Yongzheng (1722–1735), and Qianlong (1735–1796)) sourced from the developer of an ancient seal app [^4] https://www.daizigege.com/seal/. The primary reason for selecting ancient Chinese seals is that both ancient Japanese and Chinese seals are mainly inscribed with kanji characters.

### v2

@@ Abstract
Kuzushiji, a pre-modern Japanese cursive script, can currently be read and understood by only a few thousand trained experts in Japan. With the rapid development of deep learning, researchers have begun applying Optical Character Recognition (OCR) techniques to transcribe Kuzushiji into modern Japanese. Although existing OCR methods perform well on clean pre-modern Japanese documents written in Kuzushiji, they often fail to consider various types of noise, such as document degradation and seals, which significantly affect recognition accuracy. To the best of our knowledge, no existing dataset specifically addresses these challenges. To address this gap, we introduce the Degraded Kuzushiji Documents with Seals (DKDS) dataset as a new benchmark for related tasks. We describe the dataset construction process, which required the assistance of a trained Kuzushiji expert, and define two benchmark tracks: (1) text and seal detection and (2) document binarization. For the text and seal detection track, we provide baseline results using several recent versions of the You Only Look Once (YOLO) models for detecting Kuzushiji characters and seals. For the document binarization track, we present baseline results from traditional binarization algorithms, traditional algorithms combined with K-means clustering, two state-of-the-art (SOTA) Generative Adversarial Network (GAN) methods, as well as our Conditional GAN (cGAN) baseline. The DKDS dataset and the implementation code for baseline methods are available at https://ruiyangju.github.io/DKDS.
@@
Kuzushiji, Pre-modern Japanese Documents, Benchmark Dataset, Object Detection, Document Binarization, Seals
@@ 1 Introduction
Pre-modern Japanese documents were mainly written in Kuzushiji111くずし字, a form of Japanese cursive script (草書体)., which is a cursive script widely used in official records, personal letters, and literary works. Kuzushiji is characterized by complex, flowing strokes and highly variable forms, which differ markedly from modern Japanese characters [ueki2021survey]. Educational reforms in the early 20th century introduced standardized textbooks and led to reduced instruction in Kuzushiji. Following the adoption of modern kana orthography in 1946, most contemporary native Japanese speakers are unable to read Kuzushiji texts without specialized training [hashimoto2017Kuzushiji].
With the remarkable advances in deep learning for computer vision, researchers have increasingly applied neural network models to Kuzushiji character recognition. The primary datasets for Kuzushiji character recognition include Kuzushiji-MNIST, Kuzushiji-49, and Kuzushiji-Kanji [^2] Kanji refers to Chinese-origin characters used in Japanese. [clanuwat2018deep]. The Kaggle competitions organized using these datasets have further advanced both research and practical applications [kitamoto2019progress, kitamoto2020kaggle]. Together, these datasets and competitions cover a wide range of Kuzushiji characters and provide a standardized evaluation platform for researchers. With these resources, researchers can perform not only Kuzushiji character classification but also related tasks such as Kuzushiji detection and generation [soan2023], thereby advancing the development of Kuzushiji character recognition and enhancing model performance.
@@
Meanwhile, the Center for Open Data in the Humanities (CODH) has developed several Kuzushiji OCR systems that address both character-level and page-level recognition scenarios. In 2021, CODH introduced the mobile application “miwo” [clanuwat2021miwo], which enables users to capture or upload images of Kuzushiji documents and automatically transcribe them into modern Japanese. CODH also released KuroNet [clanuwat2019kuronet, lamb2020kuronet], an end-to-end OCR model based on a Residual U-Net architecture [ronneberger2015u]. Designed to capture long-range context, process extensive vocabularies, and accommodate non-standard character layouts, KuroNet supports robust recognition of pre-modern Japanese texts. On Kaggle’s Kuzushiji datasets [kitamoto2019progress, kitamoto2020kaggle], KuroNet achieved classification accuracies exceeding 90%. More recently, SakanaAI introduced Metom [Metom], a Vision Transformer [dosovitskiy2021vit]-based Kuzushiji classifier specialized for single-character recognition. Metom was trained on 2,703 types of Kuzushiji characters that appear at least five times in the large-scale Kuzushiji document provided by CODH [genjimonogatari], offering a powerful and extensible foundation for character-level analysis.
@@
Seals, serving as symbols of ownership and collector identity, are widely used in ancient and pre-modern Asian documents [li2021character]. In these documents, seals usually appear as red marks inscribed with stylized ancient scripts, and may contain the collector’s name as well as phrases reflecting personal wishes, interests, or social status. Among the various types of noise, the overlap between Kuzushiji characters and seals poses a major challenge for text and seal detection as well as document binarization.
To address this challenge, we introduce the Degraded Kuzushiji Documents with Seals (DKDS) dataset, developed in collaboration with a trained Kuzushiji expert. This novel benchmark dataset comprises degraded pre-modern Japanese documents written in Kuzushiji, with various types of seals randomly added to simulate common types of interference. As shown in Fig. 2, we define two task tracks for the DKDS dataset: (1) text and seal detection, and (2) document binarization. The former track aims to accurately detect text (i.e., Kuzushiji characters) and seal regions within the documents, while the latter focuses on preserving the primary textual content while removing noise, including stains and seals, thereby providing clean input for the subsequent OCR task.
@@
    We introduce a novel benchmark dataset, named Degraded Kuzushiji Documents with Seals (DKDS), developed in collaboration with a trained Kuzushiji expert. To the best of our knowledge, this is the first publicly available dataset specifically addressing the challenge of Kuzushiji characters overlapping with seals.
@@
    We define the text and seal detection task, which serves as a preliminary step for recognizing Kuzushiji characters and transcribing them into modern Japanese. Baseline results for this task are also presented using several recent versions of the YOLO models.
@@
    We define the document binarization task specifically for pre-modern Japanese document images. Compared to existing binarization datasets, the DKDS dataset presents additional challenges due to the overlap of Kuzushiji characters with seals. Baseline results for this task are provided, including traditional binarization algorithms, traditional algorithms combined with K-means clustering, and two state-of-the-art (SOTA) Generative Adversarial Network (GAN) methods, as well as our Conditional GAN (cGAN) baseline.
@@ 2.1 Related Dataset
The datasets relevant to this work can be categorized into three domains: (1) Kuzushiji document datasets, (2) degraded document binarization datasets, and (3) seal datasets.
Existing research [clanuwat2018deep] on pre-modern Japanese documents written in Kuzushiji lacks a dataset that specifically addresses the challenge of overlaps between characters and seals, which can significantly reduce recognition accuracy.
Most publicly available datasets for degraded document binarization focus on documents written in Latin-based alphabetic systems. For instance, the documents in the Document Image Binarization Contest (DIBCO) series datasets [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2012icfhr, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019] are primarily written in English, while the Bickley Diary (BD) [deng2010binarizationshop], Synchromedia Multispectral Ancient Document Images (SMADI) [hedjam2013historical], and Persian Heritage Image Binarization Dataset (PHIBD) [ayatollahi2013persian] all contain documents written in alphabetic scripts. In contrast, no publicly available document binarization datasets exist for pre-modern Japanese documents, particularly those written in Kuzushiji.
Furthermore, no existing degraded document datasets include seals as the primary source of interference. Yang et al. [yang2023docdiff] synthesized paired data by overlaying seal images onto clean documents; however, their synthesized documents that primarily contain modern Chinese characters, which differ significantly from the pre-modern Japanese Kuzushiji characters targeted in this work. Following a similar strategy, we also simulate seal-induced degradation by randomly adding seal images to collected pre-modern Japanese documents. Regarding the source of seal images, existing publicly available seal datasets are unsuitable for this purpose, as they either do not contain ancient Asian seals, suffer from poor image quality, or exhibit complex backgrounds. For instance, MiikeMineStamps [buitrago2021miikeminestamps] includes some Japanese seal images; however, they are generally low-resolution and feature non-uniform backgrounds, making background removal difficult.
@@ 2.2 Text and Seal Detection
The diverse shapes of ancient seals, their curved inscriptions, and the frequent occurrence of text overlap pose significant challenges for seal processing as well as text analysis and recognition. As a critical preliminary step in Kuzushiji character recognition, text detection is particularly important. For instance, DBNet [liao2020real], a widely used text detection method in OCR systems, achieved an F-measure of 85.4% at 26 frames per second on the ICDAR 2015 [karatzas2015icdar] dataset, demonstrating real-time performance with high detection accuracy.
@@
Similarly, seal detection is important for subsequent seal processing. However, the scarcity of ancient seal samples and the prevalence of low-quality or degraded images severely limit available training data and annotations for related tasks [li2021character]. Micenkov et al. [micenkov2011stamp] proposed an automatic seal segmentation system for insurance documents. When evaluated on the collected dataset of 400 document images, named StaVer, their system achieved a recall of 83% and a precision of 84%. Yu et al. [yu2023icdar] developed a modern seal dataset for official and financial contexts, introducing two tasks: seal title text detection and end-to-end seal title recognition. To the best of our knowledge, no dataset currently exists for text and seal detection in degraded pre-modern Japanese documents written in Kuzushiji.
@@
The YOLO series of models [jocher2023yolov8, wang2024yolov9, wang2024yolov10, jocher2024yolo11] achieve an excellent balance between detection accuracy and model size, demonstrating outstanding performance across a wide range of object detection tasks. Researchers have explored the application of the YOLO series of models in text detection and seal detection. Nevertheless, due to the lack of ancient seal data from pre-modern Japanese documents, the recent versions of YOLO models cannot be directly applied to this specific scenario.
2.3 Document Binarization
Document binarization is an important preprocessing step for OCR, as it effectively removes various types of noise, such as paper yellowing, text fading, ink bleeding, stains, and seals. Early work in this area primarily relied on threshold-based methods. Specifically, Otsu [otsu1979threshold] introduced the global thresholding algorithm in 1979. Subsequently, Niblack [niblack1985introduction] and Sauvola [sauvola2000adaptive] proposed local adaptive thresholding techniques in 1985 and 2000, respectively, enabling more robust binarization of documents with non-uniform illumination and complex background variations.
To facilitate systematic evaluation, the Document Image Binarization Contest (DIBCO), one of the most popular competitions in this field, released benchmark datasets from 2009 to 2019 [gatos2009icdar, pratikakis2010h, pratikakis2011icdar2011, pratikakis2012icfhr, pratikakis2013icdar, ntirogiannis2014icfhr2014, pratikakis2016icfhr2016, pratikakis2017icdar2017, pratikakis2018icdar2018, pratikakis2019icdar2019]. These datasets have played an important role in advancing both traditional and deep learning-based binarization methods by providing standardized document image evaluation benchmarks.
The introduction of GANs [goodfellow2020generative] has enabled the generation of high-quality binary document images. Suh et al. [suh2022two] proposed a two-stage GAN framework that employed six enhanced CycleGANs [zhu2017unpaired] to produce binarized document images. Building on this line of work, Ju et al. [ju2023ccdwt, ju2024three] further introduced a three-stage GAN architecture, was also composed of six improved CycleGAN modules, to advance document binarization performance. Both methods consistently achieved SOTA results on the DIBCO benchmarks.
However, neither traditional binarization methods nor SOTA GAN-based methods have been evaluated on degraded pre-modern Japanese documents written in Kuzushiji, particularly those containing seals, overlapping text, and complex degradations. This gap highlights the need to develop datasets and models that are tailored to the unique characteristics of pre-modern Japanese documents.
@@ 3.1 Raw Data Collection
The raw Kuzushiji document data were obtained from the CODH [genjimonogatari]. Specifically, we selected the dataset corresponding to book ID “200003803”, titled Genji Monogatari (The Tale of Genji) [^3] https://codh.rois.ac.jp/char-shape/book/200003803/. Examples from this book are shown in Fig. 3. This dataset was chosen for two main reasons: (1) its historical significance and widespread popularity; and (2) the availability of detailed OCR annotations provided by CODH, including 237 unique character types and a total of 11,132 characters. These data can also be used for training and evaluating OCR models in future research.
Furthermore, to date, there is no publicly available high-resolution dataset of ancient Japanese seals. To address this gap, we collected 219 seals belonging to three Chinese ancient emperors of the Qing dynasty (i.e., Kangxi (1661–1722), Yongzheng (1722–1735), and Qianlong (1735–1796)) sourced from the developer of an ancient seal app [^4] https://www.daizigege.com/seal/. The primary reason for selecting ancient Chinese seals is that both ancient Japanese and Chinese seals are mainly inscribed with Kanji characters, and no high-resolution images of ancient Japanese seals are currently publicly available.
@@
Figure 4: Examples of our collected imperial seals from the Qing dynasty, with backgrounds removed, which are used to simulate seal interference.
3.2 Construction Workflow
We constructed our dataset by producing color document images and their corresponding binarized ground-truth images from the raw data, and the overall workflow is shown in Fig. 5. To create the final document images, we combined 128 background-removed seal images with 50 Kuzushiji document images. Specifically, 40 Kuzushiji documents were selected for training, and the remaining 10 for testing. As detailed in Table 1, the 128 seal images were divided into 100 for the 40 training Kuzushiji documents, and 28 for the 10 test Kuzushiji documents. To increase the cases of overlap between seals and Kuzushiji characters in the training set, we duplicated the 100 training seal images to create 200 seals, which were then randomly added to the 40 training Kuzushiji documents. As a result, each Kuzushiji document image in the training set contains approximately 4–6 seals. In addition, we constructed easy-level (Testing-E) and difficult-level (Testing-D) versions of the test set according to the number of seals per document. Each document image in Testing-E contains approximately 2–3 seals, while each document image in Testing-D contains 14 seals.
To generate the binarization ground-truth, we employed a combined approach of neural network generation followed by manual verification and correction. As shown in Fig. 5, we first applied a pre-trained binarization model trained on the DIBCO benchmarks to binarize the raw Kuzushiji document images (without added seals), producing the initial binarization ground-truth.
Subsequently, a trained Kuzushiji expert manually verified these initial results. Even in the absence of seal interference, the initial binarization often contained errors. As shown in Fig. 6, the expert identified issues such as missing marginal notes, voiced sound symbols, periods, and unremoved stains. These errors were then manually corrected to create the final binarization ground-truth. Notably, this verification step was critical and required a trained Kuzushiji expert to ensure the validity of the process.
Figure 5: The overall workflow of the proposed DKDS dataset construction includes raw data collection, text and seal detection annotations, initial binarization ground-truth generation, verification, and manual correction. The initial binarization ground-truth was generated using a pre-trained binarization model trained on the DIBCO benchmarks, while the verification was conducted by a trained Kuzushiji expert.
Figure 6: Examples of verification and manual correction in the dataset construction: “Raw” refers to the input images, “Initial” denotes the initial binarized images produced by a pre-trained binarization model, and “Corrected” represents the final ground-truth images after verification and manual correction. The trained Kuzushiji expert performed verification to determine which elements should be preserved (e.g., marginal notes, voiced sound symbols, and periods) and which should be removed (e.g., stains).